OpenStudy (anonymous):

How do you solve (x^2 +3)^2 - (x^2 +3) - 7=0

OpenStudy (anonymous):

Substitute (x²+3) for p: p²-p-7=0

OpenStudy (anonymous):

I get these solutions for x: x= \[\sqrt{0.5(\sqrt{29}-5)}\] or x= \[-\sqrt{0.5(\sqrt{29}-5)}\]